OpenStudy (anonymous):

A homeowner wants to fence a rectangular garden using 60ft. of fencing. An existing stone wall will be used as one side of the rectangle. Find the dimensions for which the area is maximum

OpenStudy (zzr0ck3r):

so we will have 3 sides, 2w+l=60 would be the perimeter and l*w = A so l = 60-2w plug this into A A = (60-2w)*w = -2w^2+60w differentiate A with respect to W -4w+60 set equal to 0 -4w+60 = 0 w = 15 which gives L = 60-2*15 = 30 w = 15, l = 30

OpenStudy (zzr0ck3r):

we know this is a max because our function was a downward opening parabola.

OpenStudy (zzr0ck3r):

understand?

OpenStudy (anonymous):

yesss thank you very much i really apreciate it :)

OpenStudy (zzr0ck3r):

np